OpenStudy (anonymous):

find all numbers for which the rational expression is undefined. (q^3-6q)/(q^2-16)

jimthompson5910 (jim_thompson5910):

It's undefined when the denominator is zero q^2-16=0 (q+4)(q-4)=0 q+4=0 or q-4=0 q=-4 or q=4 So the expression is undefined when q=-4 or q=4

OpenStudy (amistre64):

yep, the denom goes bad when q^2 = 16; so avoid 4 and -4